The mayor of a town has proposed a plan for the annexation of an adjoining bridge. A political study took a sample of 1500 voters in the town and found that 47% of the residents favored annexation. Using the data, a political strategist wants to test the claim that the percentage of residents who favor annexation is above 44%. Determine the P-value of the test statistic. Round your answer to four decimal places.
Answer #1

Soja We 1500, § = 0.47 = * , *2705 Nul and alternate Hypothesis | Hp: P < 0:44 HI: P. 30:44 Test statistic: z= P-Po = 0.47-0.